# Online Matroid Embeddings

### v1

@@ Abstract
We introduce the notion of an online matroid embedding, which is an algorithm for mapping an unknown matroid that is revealed in an online fashion to a larger-but-known matroid. The existence of such embedding enables a reduction from the version of the matroid secretary problem where the matroid is unknown to the version where the matroid is known in advance. We show that online matroid embeddings exist for binary (and hence graphic) and laminar matroids. We also show a negative result showing that no online matroid embedding exists for the class of all matroids. Finally, we define the notion of an approximate matroid embedding, generalizing the notion of $\alpha$-partition property, and provide an upper bound on the approximability of binary matroids by a partition matroid, matching the lower bound of Dughmi et al.
@@ 1 Introduction
A common setup in online algorithms is to have a matroid whose structure is revealed to the algorithm one element at a time. The algorithm processes the elements of the ground set in sequence and at each point in time, it has access to the dependencies between the elements that have already arrived. Typical examples are the matroid secretary problem and matroid prophet inequalities.
In such problems, is there any advantage in knowing the matroid structure in advance? Imagine the following situation: we are processing an unknown matroid M, however, we know a fixed (potentially very large) matroid BigM that has an isomorphic copy of every possible matroid M and moreover, we can construct this embedding online. We will show that if such object exists, then we can reduce the version of the problem where the matroid is revealed online to the version of the problem where the matroid structure is known, by assuming our matroid is BigM.
Moreover, if the on-the-fly embedding maintains uniform random order, then the existence of such an online embedding implies that in that class the matroid secretary problem with unknown structure is no harder than the matroid secretary problem with known structure.
1.1 Two versions of the Matroid Secretary Problem
The Matroid Secretary Problem (MSP) is one of the most exciting open questions in online algorithms and a lot of effort was spent over time to understand many of its variants and special cases. In the MSP a sequence of weighted elements is presented to an algorithm in random order. The goal of the algorithm is to select elements that form an independent set of a certain matroid M and whose combined weight is in expectation an $\alpha$-approximation to the weight of the optimal basis.
In the original definition of the problem by Babaioff et al [5, 6], the algorithm has access to the matroid M in advance. The only information missing is the weight of the elements. The algorithm is presented with the labels of elements in M in random order together with their weight. Certain algorithms like Korula-Pál [22] for graphic matroids and Dinitz-Kortsarz [11] for regular matroids, rely on knowing the structure to pre-compute a decomposition of the matroid which is then used by the online algorithm.
This is in contrast with the $O(\sqrt{\log\textsf{rank}})$-approximation algorithm by Chakraborty-Lachish [9], and the more recent $O(\log\log\textsf{rank})$-approximation algorithms by Lachish [24] and Feldman-Svensson-Zenklusen [15] that only require an independence oracle over the set of elements observed so far.
The main application of the techniques we develop is to study reductions between the two versions of this problem, the known-matroid MSP and the online-revealed-matroid MSP. Before we go on and describe our main contribution, we leave the reader with two open versions of the MSP (one of each type):
MSP for binary matroids and known vectors:
An input of this problem consists of pairs $(u_{i},w_{i})$ where $u_{i}$ is a vector in $\mathbb{F}_{2}^{n}$ and $w_{i}$ is a non-negative weight. The input is presented to the algorithm in random order. The algorithm sees both the vectors and the weights and selects $\mathbb{F}_{2}$-linearly independent vectors.
MSP for graphic matroids and unknown graph:
An input consists of a graph with edge set $E$ and weights on the edges. The input is again presented to the algorithm in random order but the algorithm can only see the labels of the edges ($\textsf{edge}_{1}$, $\textsf{edge}_{2}$, $\textsf{edge}_{3}$, …) and doesn’t know the vertices they connect. Instead, the algorithm has access to an oracle which can answer for a subset of edges if it contains a cycle or not. The algorithm must select a set of edges that are independent.
An interesting consequence of our reduction in particular shows that a solution to the first problem implies a solution with the same approximation ratio for the second problem.
@@ 1.2 Our Contribution
1.2.1 Online Matroid Embedding
@@
The use of embeddings in algorithm design is an idea that has been successfully explored in other contexts, most notably, metric embeddings both in classic settings [8, 25, 3] as well as more recently in online settings [18, 4, 28]. While our motivation and main application is the matroid secretary problem, we believe that understanding maps between matroids preserving structure is an important mathematical question in its own right that can enable other algorithmic applications beyond the MSP.
1.2.2 Consequence for the MSP
@@
In Theorem 5.1 we show that the existence of an OME for a class of matroids enables a reduction from the online-revealed-matroid MSP to the known-matroid MSP. The challenge in proving this is to show that the online embedding can be used in a way that (almost) maintains uniform random arrival order. We do so by interleaving elements in BigM that are in the image of M with the remaining elements in BigM. This interleaving does not ensure uniformity of the arrival order of the elements in BigM, but it’s close to being so in total variation distance. We complete the reduction with a coupling argument that shows that the existence of an $\alpha$-competitive algorithm for the known-matroid MSP implies the existence of a $(\alpha-\epsilon)$-approximate algorithm for the online-revealed MSP.
@@
Our reduction implies that an algorithm for the MSP over binary matroids cannot meaningfully use any advance information about the matroid. This is in contrast to all known $O(1)$-competitive algorithms for special cases of binary matroids [22, 11].
1.2.3 Constructing OMEs
@@
In Section 4 (Theorem 4.2 and Theorem 4.4) we provide a complete analysis for binary matroids. Namely, for the class of binary matroids M with $n$ elements, there is an online matroid embedding into BigM the complete binary matroid $\mathbb{F}_{2}^{n}$. We also develop a technique for making the OME order-independent: we use properties of the automorphism group of $\mathbb{F}_{2}^{n}$ together with randomization to ensure that the images of the elements in M are not correlated with the arrival order. This technique is in fact more general (Theorem 4.6) and can be applied whenever the group of automorphisms of the host matroid is “sufficiently rich” in a sense that the theorem statement makes precise.
@@
We refer to online matroid embeddings where both M and BigM are of the same class as “within-class” OMEs. In Section 6 we explore whether such “within-class” OMEs can exist for graphic matroids. We show that such embeddings cannot exist, in fact we show that graphic matroids cannot be embedded in an online-fashion to regular matroids. To rule out the existence of such an online embedding, we show that if it would exist, then BigM must contain an isomorphic copy of $\mathbb{F}_{2}^{n}$. However, $\mathbb{F}^{2}_{n}$ contains an isomorphic copy of the Fano plane which is not representable over $\mathbb{F}_{3}$ [31]. Hence BigM can’t be regular.
@@
In Section 7 we give another example of an OME, namely for laminar matroids. We show (in Theorem 7.3) how to embed the class of laminar matroids M with at most $n$ elements into BigM which is a complete linear matroid of rank $n$ over any field with sufficiently many elements.
Finally in Section 8 we show an impossibility result of constructing an OME for the class of all matroids. This is shown by studying finite projective planes and showing that for those matroids, elements that haven’t arrived yet impose non-trivial constraints on the already arrived elements. As a corollary we obtain an impossibility of constructing an OME for the class of all matroids representable over fields of characteristic at least $7$.
1.2.4 Approximate OMEs
In Section 9 we extend the notion of an OME to allow distortion, i.e., the map approximately preserves the rank. We observe that the $\alpha$-partition property in Abdolazimi et al. [1] and Dughmi et al. [12] can be viewed as an approximate matroid embedding into the free matroid. Combining our formalism with their lower bounds on embedding into the free matroid, we also provide a lower bound on the distortion needed to embed the complete binary matroid into a graphic matroid (Corollary 9.3). This result is an example of the power of the formalism: with the right definitions, extending the lower bound to larger classes becomes a simple corollary. In Theorem 9.6 we show the tightness of the $\Omega(n/\log n)$ lower bound in [12] of the distortion of embedding the complete binary matroid into the free matroid by constructing an embedding achieving this distortion. Finally, we state two open questions on whether we can use distortion to bypass the impossibility results in ealier sections.
@@ 1.3 Related Work
Very recently, [30] gave a $O(1)$-competitive algorithm for the matroid secretary problem in the random assignment model when the matroid structure is not known in advance, and instead is only revealed over time.
@@
The matroid prophet inequality problem was first studied in [17]. An asymptotically optimal $(1-o(1))$-competitive algorithm for $k$-uniform matroids was given in [2]. A tight $O(1)$-competitive algorithm for the matroid prophet inequality problem was given in [23], also see [10] for the problem of maximizing submodular functions subject to matroid constraints. Constant-factor competitive algorithms can also be obtained via online contention resolution schemes (OCRS) [16]. Random-order versions of the matroid prophet inequality problem are studied in [13].
To the best of our knowledge, all these algorithms exploit that the matroid structure is known in advance. An additional difficulty for reductions of the type we present in this paper, is that typically these algorithms need to know the identity of the distribution that a certain element’s weight is drawn from. For the i.i.d. case this is obviously not an obstacle, and so our reductions apply.
@@ 2 Matroids, Morphisms, and $K$-representations
[…]
@@
Given a matroid M and an integer $k$ we will define the matroid $\textbf{M}_{[k]}$ by creating $k$ copies of each element of M. Formally, the ground set of $\textbf{M}_{[k]}$ is $\{(u,j);u\in\textbf{M},j\in[k]\}$. The rank function of $\textbf{M}_{[k]}$ is induced by the projection $\phi:\textbf{M}_{[k]}\rightarrow\textbf{M}$ that maps $(u,j)\mapsto u$, i.e., $\textsf{rank}_{\textbf{M}_{[k]}}(S)=\textsf{rank}_{\textbf{M}}(\phi(S))$. By definition, the projection $\phi$ is a matroid morphism from $_{[k]}\rightarrow\textbf{M}$ M.
@@
[…]
@@
For example, the matroids in Figure 1 can be represented by the vectors $(1,-1,0,0)$, $(0,1,-1,0)$, $(0,0,1,-1)$. As it is the case for graphic matroids, the matroid description has no information about vectors and the representation is again not unique. An equally good representation is $(1,0,0),(0,1,0),(0,0,1)$.
@@
We will denote by $\textbf{U}_{n,r}$ the uniform matroid of $n$ elements and rank $r$. This is the matroid with ground set $[n]$ and whose rank function is $\textsf{rank}_{\textbf{U}_{n,r}}(S)=\min(r,|S|)$. We call the $\textbf{Fr}_{n}:=\textbf{U}_{n,n}$ the free matroid of rank $n$, i.e., a matroid of $n$ elements in which every element is independent.

### v2

@@ Abstract
We introduce the notion of an online matroid embedding, which is an algorithm for mapping an unknown matroid that is revealed in an online fashion to a larger-but-known matroid. We establish the existence of such an embedding for binary matroids, and use it to relate variants of the binary matroid secretary problem to each other, showing that seemingly simpler problems are in fact equivalent to seemingly harder ones (up to constant-factors). Specifically, we show this to be the case for the version of the matroid secretary problem in which the matroid is not known in advance, and where it is known in advance. We also show that the version with known matroid structure, is equivalent to the problem where weights are not fully adversarial but drawn from a known pairwise-independent distribution.
@@ 1 Introduction
A common setup in online algorithms is to have a matroid whose structure is revealed to the algorithm one element at a time. The algorithm processes the elements of the ground set in sequence, and at each point in time, it has access to the dependencies between the elements that have already arrived. Typical examples include the famous matroid secretary problem (MSP) [8, 9, 32, 22] and matroid prophet inequalities [13, 31].
In such problems, is there any advantage in knowing the matroid structure in advance? Imagine the following situation: we are processing an unknown matroid M; however, we know a fixed (potentially very large) matroid BigM that has an isomorphic copy of every possible matroid M, and we can construct this embedding online. We will show that if such an object exists, then we can reduce the version of the problem where the matroid is revealed online to the version of the problem where the matroid structure is known, by assuming our matroid is BigM. Moreover, if the on-the-fly embedding maintains uniform random order, then the existence of such an online embedding implies that in that class, the matroid secretary problem with unknown structure is no harder than the matroid secretary problem with known structure.
Important recent progress on the MSP has established that it is equivalent to the matroid prophet secretary problem with correlated distributions [19, 18]. Though seemingly unrelated, another consequence of the existence of such online embeddings will be that—for certain matroids— this equivalence holds even if we impose pairwise independence.
@@ 1.1 Our Contribution
Online Matroid Embedding
@@
The use of embeddings in algorithm design is an idea that has been successfully explored in other contexts, most notably, metric embeddings both in classic settings [11, 33, 6] as well as more recently in online settings [25, 7, 37]. While our motivation and main application is the matroid secretary problem, we believe that understanding maps between matroids preserving structure is an important mathematical question in its own right that can enable other algorithmic applications beyond the matroid secretary problem.
Consequences for the MSP
We use the concept of online matroid embeddings to gain insights into the complexity of the matroid secretary problem (MSP) on binary matroids (see Section 2). Specifically, we relate different variants of the problem to each other and show that seemingly simpler ones are actually equivalent to harder ones (up to constant factors). See Figure 1 for an overview of the reductions that we establish in this paper.
The three variants we are interested in are: (1) the online-revealed-matroid MSP, where the matroid is a priori unknown to the online algorithm and the algorithm has access to an independence oracle on the already arrived elements, (2) the known-matroid MSP, where the structure of the matroid is known to the algorithm in advance, and (3) the prophet MSP, where the matroid structure is known in advance and additionally the weights of the elements are drawn from a known, but possibly correlated distribution.
known-matroid
MSP
prophet MSP
w/ pairwise independence
online-revealed-matroid
MSP
[…]
Figure 1: Reductions for binary matroids. We use $P\succeq Q$ to indicate that $P$ is harder than $Q$, and we use $\succeq_{\epsilon}$ and $\succeq_{C}$ to designate an additive $\varepsilon$ or multiplicative factor $C$ loss in approximation.
Clearly, the online-revealed-matroid MSP is harder than the known-matroid MSP and the known-matroid MSP is harder than the prophet MSP, in the sense that if we have an $\alpha$-approximation for one problem, then we also have an $\alpha$-approximation for the other. Two main implications of our work are “inverses” of these statements for binary matroids, that hold up to a constant-factor loss, and apply even if we impose pairwise-independence in the prophet MSP. Such reductions between different average-case problems are notoriously difficult to achieve, as they need to ensure or maintain rather stringent assumptions on the input distribution that are essential for the target algorithm to be applicable in a meaningful way, and the required properties are easily disrupted.
@@
Step 1: A Reduction From Online-Revealed Matroid MSP to Known-Matroid MSP. In Theorem 5.1 we show that the existence of an OME for a class of matroids enables a reduction from the online-revealed-matroid MSP to the known-matroid MSP. The challenge in proving this is to show that the online embedding can be used in a way that (almost) maintains uniform random arrival order. More precisely, let M be the unknown matroid that is revealed to the algorithm in an online fashion and let $f$ be an OME into BigM. Now consider the reduction: upon arrival of an element-weight pair $(e,w_{e})$ in matroid M at iteration $t$, we construct the corresponding element-weight pair $(f(e),w_{e})$ as an input to MSP on BigM at iteration $t$. However, this is not a valid input to known-matroid MSP on BigM as it does not construct a random arrival order over the elements that are not in the image of M.
To overcome the shortcoming of the above simple reduction, we interleave the elements in BigM that are in the image of M with the remaining elements in BigM. In the proof of Theorem 5.1, our main technical argument shows that, while this interleaving does not ensure uniformity of the arrival order of the elements in BigM, it leads to an arrival order over elements in BigM that is close to uniformly random arrival order in total variation distance (Section 5.1). We then complete the reduction with a coupling argument that shows that the existence of an $\alpha$-competitive algorithm for the known-matroid MSP implies the existence of an $(\alpha-\epsilon)$-competitive algorithm for the online-revealed-matroid MSP.
@@
Together with the existence of an OME that maps binary matroids into the complete binary matroid (see Section 4 and discussion below), our reduction implies that an algorithm for the MSP over binary matroids cannot meaningfully use any advance information about the matroid (Corollary 5.2). This is in contrast to all known $O(1)$-competitive algorithms for special cases of binary matroids [30, 16].
Step 2: A Reduction from Prophet MSP w/ Pairwise Independence to Known-Matroid MSP. In Theorem 5.3, we show that the existence of an OME from a class of matroids $\mathcal{C}$ to BigM satisfying a $2$-transitivity property (that is satisfied by complete binary matroids, see definition in Section 2), allows to translate an $\alpha$-competitive algorithm for prophet MSP with pairwise-independent weight distribution on BigM into a $C\cdot(\alpha-o(1))$-competitive algorithm for known-matroid MSP on matroid $\textbf{M}\in\mathcal{C}$, for some constant $C>0$.
To establish Theorem 5.3, we build on [19, 18] and show how to reduce prophet MSP with arbitrary correlation on $\textbf{M}\in\mathcal{C}$ to prophet MSP with pairwise-independent weight distribution on BigM. [^1] A technical detail that we are ignoring here is that our reduction is from a restricted version of the prophet MSP with arbitrary correlation, which results in an additional constant-factor loss. To prove this, we first show that any uniformly random automorphism $f:\textbf{BigM}\rightarrow\textbf{BigM}$ satisfies the following property: for any pair of elements $e,e^{\prime}$, $\Pr[f(e)=e^{\prime}]=\frac{1}{n}$ and for any two pairs of independent elements $e_{1},e_{2}$ and $e_{1}^{\prime},e_{2}^{\prime}$, $\Pr[f(e_{1})=e_{1}^{\prime}\land f(e_{2})=e_{2}^{\prime}]=\frac{1}{n\cdot(n-1)}$, where $n=|\textbf{BigM}|$ (Lemma 3.2). This property allows us to construct an “almost pairwise independent” randomized OME $f^{\prime}:\textbf{M}\rightarrow\textbf{BigM}$ by simply composing the given OME with a uniformly random automorphism on BigM. The resulting weight distribution is approximately pairwise independent in the sense that for any pair of elements $e,e^{\prime}\in\textbf{BigM}$ and weights $w,w^{\prime}$ it holds that $|\Pr[w(e)=w]\cdot\Pr[w(e^{\prime})=w^{\prime}]-\Pr[w(e)=w^{\prime}\land w(e^{\prime})=w^{\prime}]|=O\left(\frac{1}{n^{2}}\right)$.
To conclude the proof, we show that there exists a pairwise-independent distribution that is close to the induced weight distribution over BigM (Theorem 5.9). This is one of the most technical results of the paper (see discussion below, and Section B). Combining Theorem 5.9 with a coupling argument similar to the one in our other reduction completes the proof.
We note that constructing an exactly $k$-wise independent distribution from an approximately $k$-wise independent distribution has been studied in previous work [2, 4, 1]. However, their techniques focus on a set of Bernoulli random variables with identical marginals [2, 1] or “uniformity” of the underlying random variables [4]—both conditions do not hold in our setting as the weight distribution in the prophet MSP instance can be arbitrarily correlated. In fact, in both works [2, 4], they show that if the random variables $X_{1},\dots,X_{n}$ satisfy “uniformity” and $|\mathbb{E}[X_{i}\cdot X_{j}]-\mathbb{E}[X_{i}]\cdot\mathbb{E}[X_{j}]|\leq\varepsilon$ then there exists pairwise independent random variables $\tilde{X}_{1},\dots,\tilde{X}_{n}$ within a distance of $O(n^{2}\cdot\varepsilon)$ — which is not enough for our purpose as $\varepsilon=\Theta(1/n^{2})$ in our case.
To obtain Theorem 5.9, we construct an explicit pairwise-independent weight distribution over BigM by a sequence of “small” perturbations to a naturally induced “almost” pairwise-independent distribution. At each step, we perturb the original distribution such that $\omega(1)$ many pairs of random variables end up being independent (Procedure LABEL:proc:one and Procedure LABEL:proc:two in Section B) while always decreasing the pairwise correlation of the rest of the pairs (Lemma B.1). Then the main technical work is devoted to showing that the total deviation through our procedures is in the order of $\varepsilon\cdot o(n^{2})$ (Section B.3 and Section B.4), which, combined with the fact that $\varepsilon=O(1/n^{2})$ leads to the desired result. We believe that our idea of sequentially constructing small perturbations would find further applications to obtain exact pairwise (or $k$-wise) independent distributions from their approximate counterparts in other settings.
Constructing OMEs
@@
In Section 4 (Theorem 4.2 and Theorem 4.4), we provide a complete analysis for binary matroids. Namely, for the class of binary matroids M with $n$ elements, there is an online matroid embedding into BigM the complete binary matroid $\mathbb{F}_{2}^{n}$. We also develop a technique for making the OME order-independent: we use properties of the automorphism group of $\mathbb{F}_{2}^{n}$ together with randomization to ensure that the images of the elements in M are not correlated with the arrival order. This technique is in fact more general (Theorem 4.6) and can be applied whenever the group of automorphisms of the host matroid is “sufficiently rich” in a sense that the theorem statement makes precise.
@@
We refer to online matroid embeddings where both M and BigM are of the same class as “within-class” OMEs. In Section 6.1 we explore whether such “within-class” OMEs can exist for graphic matroids. We show that such embeddings cannot exist, in fact we show that graphic matroids cannot be embedded in an online-fashion to regular matroids. To rule out the existence of such an online embedding, we show that if it would exist, then BigM must contain an isomorphic copy of $\mathbb{F}_{2}^{n}$. However, $\mathbb{F}_{2}^{n}$ contains an isomorphic copy of the Fano plane which is not representable over $\mathbb{F}_{3}$ [41]. Hence BigM can’t be regular.
@@
In Section 6.2 we give another example of an OME, namely for laminar matroids. We show (in Theorem 6.6) how to embed the class of laminar matroids M with at most $n$ elements into BigM which is a complete linear matroid of rank $n$ over any field with sufficiently many elements.
Finally in Section 6.3 we show an impossibility result of constructing an OME for the class of all matroids. This is shown by studying finite projective planes and showing that for those matroids, elements that haven’t arrived yet impose non-trivial constraints on the already arrived elements. As a corollary we obtain an impossibility of constructing an OME for the class of all matroids representable over fields of characteristic at least $7$.
Approximate OMEs
In Section 7 we extend the notion of an OME to allow distortion, i.e., the map approximately preserves the rank. We observe that the $\alpha$-partition property in [3] and [17] can be viewed as an approximate matroid embedding into the free matroid.
First, combining our formalism with their lower bounds on embedding into the free matroid, we also provide a lower bound on the distortion needed to embed the complete binary matroid into a graphic matroid (Corollary 7.3). This result is an example of the power of the formalism: with the right definitions, extending the lower bound to larger classes becomes a simple corollary.
In Theorem 7.6 we show the tightness of the $\Omega(n/\log n)$ lower bound in [17] of the distortion of embedding the complete binary matroid into the free matroid by constructing an embedding achieving this distortion. We also show that there is no constant approximate online embedding of the class of graphical matroids into a free matroid when the underlying matroid is not known upfront (Theorem 7.7). Therefore, any constant competitive algorithm for unknown graphical matroid secretary that relies on constructing an online embedding of the graph into a free-matroid has to exploit the random arrival order of the underlying elements or develop new techniques that do not rely on online embedding into a free matroid.
1.2 Discussion and Significance of Results
We believe that the existence or non-existence of (approximate) online matroid embeddings can shed new light on different classes of matroids and how they relate to each other. In this work, we demonstrate two implications for the matroid secretary problem.
Our first implication (Theorem 5.1) offers the first formalization of the intuition that, in general, advance knowledge of the matroid structure should not help in the design of a constant-competitive algorithm for the matroid secretary problem. In light of this, it would be interesting to develop algorithms for classes of matroids for which constant-competitive algorithms exist when the algorithm has advance knowledge of the matroid structure [e.g., 30, 16].
Our second implication (Theorem 5.3), in turn, presents a novel “line of attack” for obtaining such an algorithm for the class of binary matroids (for which no constant-competitive algorithm is known). While it was already known that it suffices to find such an algorithm for the secretary prophet version with correlated weights [19, 18], general correlated weight distributions offer little additional structure. Our result shifts the challenge away from intractable arbitrary correlations, towards the better-understood realm of pairwise independent distributions. Pairwise independent distributions admit powerful tools like concentration inequalities and have found application in areas such as hashing and constructions of pseudo-random generators (for more details, see surveys [35, 42]), as well as prophet inequalities [12].
@@ 1.3 Related Work
Very recently, [40] gave a $O(1)$-competitive algorithm for the matroid secretary problem in the random assignment model when the matroid structure is not known in advance, and instead is only revealed over time. In a similar spirit, [39] presents an online contention resolution scheme for graphic matroids, that uses almost no advance information about the graph. However, they assume that the endpoints of the edges are revealed upon their arrival which leads to an obvious OME into a graphical matroid.
@@
The matroid prophet inequality problem was first studied in [24]. An asymptotically optimal (1o(1))(1-o(1))-competitive algorithm for $k$-uniform matroids was given in [5]. A tight $O(1)$-competitive algorithm for the matroid prophet inequality problem was given in [31], also see [15] for the problem of maximizing submodular functions subject to matroid constraints. Constant-factor competitive algorithms can also be obtained via online contention resolution schemes (OCRS) [23]. Random-order versions of the matroid prophet inequality problem are studied in [20].
To the best of our knowledge, all these algorithms exploit that the matroid structure is known in advance. An additional difficulty for reductions of the type we present in this paper, is that typically these algorithms need to know the identity of the distribution that a certain element’s weight is drawn from. For the i.i.d. case this is obviously not an obstacle, and so our reductions apply. We believe that extensions of our techniques might shed further light on the variant of the matroid prophet inequality problem, in which the matroid is revealed online.
@@ 2 Matroids, Morphisms, and $K$-representations
We refer to the set of automorphisms $\textbf{M}\rightarrow\textbf{M}$ as $\text{Aut}(\textbf{M})$, which forms a group under composition, i.e., given $f,g\in\text{Aut}(\textbf{M})$, then $f\circ g\in\text{Aut}(\textbf{M})$ (and $\circ$ satisfies the group axioms).
@@
Given a matroid M and an integer $k$ we will define the matroid $\textbf{M}_{[k]}$ by creating $k$ copies of each element of M. Formally, the ground set of $\textbf{M}_{[k]}$ is $\{(u,j);u\in\textbf{M},j\in[k]\}$. The rank function of $\textbf{M}_{[k]}$ is induced by the projection $\phi:\textbf{M}_{[k]}\rightarrow\textbf{M}$ that maps $(u,j)\mapsto u$, i.e., $\textsf{rank}_{\textbf{M}_{[k]}}(S)=\textsf{rank}_{\textbf{M}}(\phi(S))$. By definition, the projection $\phi$ is a matroid morphism from $\textbf{M}_{[k]}\rightarrow\textbf{M}$.
@@
Given two matroids M and N, we define their direct sum $\textbf{M}\oplus\textbf{N}$ as the matroid whose ground set is the disjoint union of the ground sets of M and N and $\textsf{rank}_{\textbf{M}\oplus\textbf{N}}(S)=\textsf{rank}_{\textbf{M}}(S\cap\textbf{M})+\textsf{rank}_{\textbf{N}}(S\cap\textbf{N})$ for all $S$ in the disjoint union of ground sets.
@@
For example, the matroids in Figure 2 can be represented by the vectors (1,1,0,0)(1,-1,0,0), (0,1,1,0)(0,1,-1,0), (0,0,1,1)(0,0,1,-1). As it is the case for graphic matroids, the matroid description has no information about vectors and the representation is again not unique. An equally good representation is $(1,0,0),(0,1,0),(0,0,1)$.
@@
We will denote by $\textbf{U}_{n,r}$ the uniform matroid of $n$ elements and rank $r$. This is the matroid with ground set $[n]$ and whose rank function is $\textsf{rank}_{\textbf{U}_{n,r}}(S)=\min(r,|S|)$. We call the $\textbf{Fr}_{n}:=\textbf{U}_{n,n}$ the free matroid of rank $n$, i.e., a matroid of $n$ elements in which every set is independent.
@@
22-transitive Matroid
We say that a simple matroid M (loop-free and no parallel elements) is $2$-transitive if for any pair of independent sets of size two, $\{e_{1},e_{1}^{\prime}\},\{e_{2},e_{2}^{\prime}\}\in\textbf{M}$ there exists an automorphism $f\in\operatorname{Aut}(\textbf{M})$ satisfying $f(e_{1})=e_{1}^{\prime}$ and $f(e_{2})=e_{2}^{\prime}$. There are several matroids that satisfies the $2$-transitive property including complete affine matroids, complete projective matroids, free matroids, and their truncations [29].